# Many-body localization crossover is sharper in quasiperiodic spin chains

### v1

@@ Acknowledgments.
Figure S.3: (a) Different choices of $\delta_{\overline{r}}$ for the near ergodic disorder strength $W_{\overline{r}}^{T}$, together with the critical point $W_{\overline{r}}^{*}$. The dotted lines represent linear scaling in $L$, while the dashed line represents a first-order polynomial in $1/L$; (b) Average gap ratio $\overline{r}$ as a function of $W$ near its COE value. The red (black) cross represents the choice of the near ergodic region for $\delta_{1}$ ($\delta_{2}$), while the filled (open) circles represent the data obtained for $\kappa_{0}$ ($\kappa_{1}$) at different system sizes.
@@
In Fig. S.3(a), we show the details on the determination of the near ergodic disorder strength $W_{\overline{r}}^{T}$ for different choices of $\delta_{\overline{r}}$, together with the critical disorder strength $W_{\overline{r}}^{*}$ (blue diamonds). Similarly to the KIM with random disorder, the behavior of $W_{\overline{r}}^{T}$ starts to deviate from the linear scaling around $L=16$ for two different choices of $\delta_{\overline{r}}$ and this may indicate the stability of the MBL phase. The points where both values were extracted are illustrated in Fig. S.3(b), and it shows that the choice of $\delta_{2}=0.03$ leads to a faster slowdown than the $\delta_{1}=0.01$ choice, consistent with the sharpness of the MBL crossover.
@@
The critical disorder strength $W_{\overline{r}}^{*}$, despite some fluctuations that are inherent from the QP potential, can be reasonably well described by a first-order polynomial in $1/L$, with the fitting errors being shown by the shaded blue area. If we perform an extrapolation of this functional form, the crossing with the linear fitting of $W_{\overline{r}}^{T}$ happens at $L\approx 26$ ($L\approx 25$) for the choice of $\delta_{1}$ ($\delta_{2}$), as shown by the cross symbols in Fig. S.3(a).
S-IV.2 Slopes
Figure S.4: (a) Derivatives of $\overline{r}(W)$ for different system sizes across the MBL crossover, where the red circles represent the biggest slope and the blue cross is the slope value at $\overline{r}_{\mathrm{avg}}\approx 0.458$; (b) A direct comparison between the exponential behavior of these two different points. Despite minor differences, both curves scale with similar exponent ($\lambda\approx 0.1$).
@@
We then obtain the slopes ($\alpha$) across the entire MBL crossover and extract their minimum values together with the corresponding disorder strength, $W_{\overline{x}}^{\alpha}$. We highlight the minimum value of $\alpha$ as the red circles in Fig. S.4(a), which clearly shows the exponential behavior of the slopes. To test the robustness of this exponential flow towards a non-analytic behavior of QP models, we also extract the slopes at the average value between $\mathrm{COE}$ and $\mathrm{Poisson}$ expectation values, $\overline{x}_{\mathrm{avg}}$. This value corresponds to $\overline{r}_{\mathrm{avg}}\approx 0.458$ if we consider average gap ratio as the quantity of interest and $\overline{s}_{\mathrm{avg}}\approx 0.5$ if we consider the rescaled entanglement entropy. In Fig. S.4(a), we show how both points approach each other as we increase the system size $L$, and this suggests the KIM with QP potential is very close to a step-like function behavior for $\overline{r}$. The exponential behavior of both points is shown in Fig. S.4(b), demonstrating the consistency of the exponential behavior, with both curves having a similar exponent ($\lambda\approx 0.093(4)$ for the maximum and $\lambda\approx 0.101(3)$ for the average point).
To reinforce there are no qualitative differences between different choices for the slopes, we show in Fig. S.5 how $\alpha$ scales if we take them at a fixed point ($\overline{x}_{\mathrm{avg}}$). The exponential behavior persists for both models with QP potential, with a similar scaling exponent of the maximum slope. For the average gap ratio, panel (a), we obtain an exponent $\lambda=0.101(3)$ for the KIM and $\lambda=0.100(4)$ for the HSC, while the rescaled entanglement entropy, panel (b), have exponents $\lambda=0.078(2)$ for the KIM and $\lambda=0.070(2)$ for the HSC. Similarly, models with random disorder exhibit a linear behavior with the system size, showing comparable scaling parameters to those observed for $|\alpha_{\mathrm{max}}|$.
Figure S.5: Sharpness of the ETH-MBL crossover in disorder spin chains taken at (a) $\overline{r}_{\mathrm{avg}}\approx 0.458$ and (b) $\overline{s}_{\mathrm{avg}}\approx 0.5$. Similarly to the maximum slope (see the main text), the slopes taken at the average values $|\alpha_{\mathrm{avg}}|$ grows exponentially $|\alpha_{\mathrm{avg}}|\propto e^{\lambda L}$ for the QP models and linearly for the RD models. The exponents $\lambda$ are also similar, with $\lambda=0.101(3)$ for the KIM and $\lambda=0.100(4)$ for the HSC, showing the robustness of the exponential scaling.
S-IV.3 Uncertainties on finite-size quantities
@@
$\sigma_{x}=\frac{(\langle(x_{S}-\overline{x})^{2}\rangle)^{1/2}}{N_{\mathrm{dis}}^{1/2}}$ (S.1)
@@
S-V Finite-size scaling analysis for QP KIM
@@
Predictions on the functional form of the correlation length are generally lacking for QP models. A few real-space renormalization group studies predict a power-law-like scaling [167], while the absence of a mechanism akin to an avalanche scenario presently does not lead to any BKT-like predictions. However, the observation that slopes grow exponentially with increasing system size for a fixed value of $\bar{x}$ motivates us to consider both functional forms of the correlation function.
@@
S-V.1 Fixed crossover disorder strength.
@@
S-V.2 Size dependent crossover disorder strength.
@@
Figure S.6: Assuming a BKT-like correlation function $\xi_{B}$ and a functional form for drifts $W_{c}(L)=W_{0}+W_{1}L^{\gamma}$ (a) and (b) shows how the logarithm of the cost-function changes in the ($\gamma$,$W_{0}$) space for $\bar{r}$ and $\bar{s}$ respectively. While (c) and (d) make a comparison of the minimized $W_{c}(L)$, both for a global (GM) and physical (PM) minima, with disorder strengths $W_{\bar{x}}^{T}$ and $W_{\bar{x}}^{\alpha}$ for $\bar{r}$ and $\bar{s}$ respectively.
In the parameter space of $\mathcal{C}_{\bar{x}}$ with a BKT-like correlation function with system size drifts in crossover disorder strengths $W_{c}(L)=W_{0}+W_{1}L^{\gamma}$ over $(W_{0},\gamma)$ as shown in Fig. S.6 we find an island with global minima (GM) that extends to $W_{0}<0$. The parameter values resulting from these minima can be non-physical, leading to a negative critical disorder strength for small values. We also find an island with physically relevant minima (PM) that has $W_{0}>0$ and $-1<\gamma<0$. The parameters $(W_{0},W_{1},\gamma,b)$ found with PM are for $\bar{r}\rightarrow(7.34,-9.31,-0.21,1.29)$ while for $\bar{s}\rightarrow(6.94,-9.82,-0.22,2.40)$. Indicating the system size-dependent drift of the crossover disorder strength saturating at $W\approx 7$.
S-V.3 Numerical details on cost-function

### v2

@@ B.1 Finite-size drifts
Figure 7: (a) Different choices of $\delta_{\overline{r}}$ for the near ergodic disorder strength $W_{\overline{r}}^{T}$, together with the critical point $W_{\overline{r}}^{*}$. The dotted lines represent linear scaling in $L$, while the dashed line represents a first-order polynomial in $1/L$; (b) Average gap ratio $\overline{r}$ as a function of $W$ near its COE value. The red (black) cross represents the choice of the near ergodic region for $\delta_{1}$ ($\delta_{2}$), while the filled (open) circles represent the data obtained for $\kappa_{0}$ ($\kappa_{1}$) at different system sizes.
@@
In Fig. 7(a), we show the details on the determination of the near ergodic disorder strength $W_{\overline{r}}^{T}$ for different choices of $\delta_{\overline{r}}$, together with the critical disorder strength $W_{\overline{r}}^{*}$ (blue diamonds). Similarly to the KIM with random disorder, the behavior of $W_{\overline{r}}^{T}$ starts to deviate from the linear scaling around $L=16$ for two different choices of $\delta_{\overline{r}}$ and this may indicate the stability of the MBL phase. The points where both values were extracted are illustrated in Fig. 7(b), and it shows that the choice of $\delta_{2}=0.03$ leads to a faster slowdown than the $\delta_{1}=0.01$ choice, consistent with the sharpness of the MBL crossover.
@@
The critical disorder strength $W_{\overline{r}}^{*}$, despite some fluctuations that are inherent from the QP potential, can be reasonably well described by a first-order polynomial in $1/L$, with the fitting errors being shown by the shaded blue area. If we perform an extrapolation of this functional form, the crossing with the linear fitting of $W_{\overline{r}}^{T}$ happens at $L\approx 26$ ($L\approx 25$) for the choice of $\delta_{1}$ ($\delta_{2}$), as shown by the cross symbols in Fig. 7(a).
B.2 Slopes
@@
We then obtain the slopes ($\alpha$) across the entire MBL crossover and extract their minimum values together with the corresponding disorder strength, $W_{\overline{x}}^{\alpha}$. We highlight the minimum value of $\alpha$ as the red circles in Fig. 2(a), which clearly shows the exponential behavior of the slopes. To test the robustness of this exponential flow towards a non-analytic behavior of QP models, we also extract the slopes at the average value between $\mathrm{COE}$ and $\mathrm{Poisson}$ expectation values, $\overline{x}_{\mathrm{avg}}$. This value corresponds to $\overline{r}_{\mathrm{avg}}\approx 0.458$ if we consider average gap ratio as the quantity of interest and $\overline{s}_{\mathrm{avg}}\approx 0.5$ if we consider the rescaled entanglement entropy. In Fig. 2(a), we show how both points approach each other as we increase the system size $L$, and this suggests the KIM with QP potential is very close to a step-like function behavior for $\overline{r}$. The exponential behavior of both points is shown in Fig. 2(b), demonstrating the consistency of the exponential behavior, with both curves having a similar exponent ($\lambda\approx 0.093(4)$ for the maximum and $\lambda\approx 0.101(3)$ for the average point).
To reinforce there are no qualitative differences between different choices for the slopes, we show in Fig. 5 how $\alpha$ scales if we take them at a fixed point ($\overline{x}_{\mathrm{avg}}$). The exponential behavior persists for both models with QP potential, with a similar scaling exponent of the maximum slope. For the average gap ratio, panel (a), we obtain an exponent $\lambda=0.101(3)$ for the KIM and $\lambda=0.100(4)$ for the HSC, while the rescaled entanglement entropy, panel (b), have exponents $\lambda=0.078(2)$ for the KIM and $\lambda=0.070(2)$ for the HSC. Similarly, models with random disorder exhibit a linear behavior with the system size, showing comparable scaling parameters to those observed for $|\alpha_{\mathrm{max}}|$.
B.3 Uncertainties on finite-size quantities
@@
$\sigma_{x}=\frac{(\langle(x_{S}-\overline{x})^{2}\rangle)^{1/2}}{N_{\mathrm{dis}}^{1/2}}$ (7)
@@
Appendix C Finite-size scaling analysis for QP KIM
@@
Predictions on the functional form of the correlation length are generally lacking for QP models. A few real-space renormalization group studies predict a power-law-like scaling [118, 119], while the absence of a mechanism akin to an avalanche scenario presently does not lead to any BKT-like predictions. However, the observation that slopes grow exponentially with increasing system size for a fixed value of $\bar{x}$ motivates us to consider both functional forms of the correlation function.
@@
C.1 Fixed crossover disorder strength.
@@
C.2 Size dependent crossover disorder strength.
@@
Figure 8: Assuming a BKT-like correlation function $\xi_{B}$ and a functional form for drifts $W_{c}(L)=W_{0}+W_{1}L^{\gamma}$ (a) and (b) shows how the logarithm of the cost-function changes in the ($\gamma$,$W_{0}$) space for $\bar{r}$ and $\bar{s}$ respectively. While (c) and (d) make a comparison of the minimized $W_{c}(L)$, both for a global (GM) and physical (PM) minima, with disorder strengths $W_{\bar{x}}^{T}$ and $W_{\bar{x}}^{\alpha}$ for $\bar{r}$ and $\bar{s}$ respectively.
In the parameter space of $\mathcal{C}_{\bar{x}}$ with a BKT-like correlation function with system size drifts in crossover disorder strengths $W_{c}(L)=W_{0}+W_{1}L^{\gamma}$ over $(W_{0},\gamma)$ as shown in Fig. 8 we find an island with global minima (GM) that extends to $W_{0}<0$. The parameter values resulting from these minima can be non-physical, leading to a negative critical disorder strength for small values. We also find an island with physically relevant minima (PM) that has $W_{0}>0$ and $-1<\gamma<0$. The parameters $(W_{0},W_{1},\gamma,b)$ found with PM are for $\bar{r}\rightarrow(7.34,-9.31,-0.21,1.29)$ while for $\bar{s}\rightarrow(6.94,-9.82,-0.22,2.40)$. Indicating the system size-dependent drift of the crossover disorder strength saturating at $W\approx 7$.
C.3 Numerical details on cost-function